OpenStudy (anonymous):

How to find the derivative for f(x)= 5x+1/x, x=-1

OpenStudy (anonymous):

you have a choice. a) you can use the annoying quotient rule \[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\] or you can write \[f(x)=5-\frac{1}{x}\] and take the derivative of that, which is much simpler

OpenStudy (anonymous):

oops last line should have been \[f(x)=5+\frac{1}{x}\]

OpenStudy (anonymous):

you go (5x+1)*x^-1 and then take derivative

OpenStudy (anonymous):

derivative of a constant is 0 derivative of \(\frac{1}{x}\) is \(-\frac{1}{x^2}\) replace x by \(-1\) to get your answer

OpenStudy (anonymous):

mathdood method would work, but i would not recommend it easier just to divide each term by x and go from there

OpenStudy (anonymous):

Okay i know to to solve it with the qoutient rule but not how to solve.it with the derivative rule

OpenStudy (anonymous):

no no don't use the quotient rule

OpenStudy (anonymous):

the quotient rule is a derivative rule, but you don't want to use it here because there is a much much simpler way

OpenStudy (anonymous):

\[f(x)=5+\frac{1}{x}\] \[f'(x)=-\frac{1}{x}^2\] \[f'(-1)=-\frac{1}{(-1)^2}=-1\]

OpenStudy (anonymous):

But the answer in my textbook is 4.....

OpenStudy (anonymous):

\[ f(x)= 5x+\frac1x\] \[ f'(x)=5-\frac{1}{x}^2 \] \[ f'(-1)=5-\frac{1}{(-1)^2}=5-1=4\]